### v4

# LookAhead Tuning: Safer Language Models via Partial Answer Previews

Abstract.
Fine-tuning enables large language models (LLMs) to adapt to specific domains, but often compromises their previously established safety alignment. To mitigate the degradation of model safety during fine-tuning, we introduce LookAhead Tuning, a lightweight and effective data-driven approach that preserves safety during fine-tuning. The method introduces two simple strategies that modify training data by previewing partial answer prefixes, thereby minimizing perturbations to the model’s initial token distributions and maintaining its built-in safety mechanisms. Comprehensive experiments demonstrate that LookAhead Tuning effectively maintains model safety without sacrificing robust performance on downstream tasks. Our findings position LookAhead Tuning [^1] https://github.com/zjunlp/LookAheadTuning. as a reliable and efficient solution for the safe and effective adaptation of LLMs.
Large Language Models, Fine-Tuning, Security
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining; February 22–26, 2026; Boise, ID, USAbooktitle: Proceedings of the Nineteenth ACM International Conference on Web Search and Data Mining (WSDM ’26), February 22–26, 2026, Boise, ID, USAdoi: 10.1145/3773966.3779370isbn: 979-8-4007-2292-9/2026/02ccs: Computing methodologies Natural language generationccs: Security and privacy Usability in security and privacyccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Natural language processing
1. Introduction
Fine-tuning effectively enhances the capabilities of large language models (LLMs), for example, fine-tuning chain-of-thought (DBLP:conf/nips/Wei0SBIXCLZ22) data can improve reasoning abilities (DBLP:journals/corr/abs-2411-14405; muennighoff2025s1), and adapting to domain-specific task data can substantially boost task performance (DBLP:journals/corr/abs-2303-18223; DBLP:conf/acl/Zhang0ZX24). However, these performance gains often come at the cost of safety alignment (yang2023shadow; lermen2023lora; zhan2023removing; yi2024vulnerability; alber2025medical). Even benign training data can induce catastrophic degradation of protective mechanisms during vanilla fine-tuning processes (qi2023fine; shallowalign; pelrine2023exploiting; he2024s). This raises a critical challenge: how can we endow models with new capabilities without sacrificing their safety? Existing work highlights several obstacles to achieving this goal, including the neglect of the impact of benign data on model safety (huang2024vaccine), limited methodological adaptability (wei2024assessing), and high computational demands (DBLP:conf/acl/YangPFWCZL24). These issues underscore the need for a solution that is both effective and resource-efficient.
Figure 1. Our method maintains safety alignment comparable to the seed model, while achieving task performance improvements equivalent to vanilla fine-tuning.
Figure 2. Overview of Training Data Modification: Vanilla Fine-Tuning; LookAhead-Tuning with True Answer ($m=6$); LookAhead-Tuning with Virtual Answer ($P=$ “Let’s solve this problem.”). Inference data is unchanged.
To address these challenges, understanding how to mitigate forgetting of the model’s safety knowledge (DBLP:journals/corr/abs-2308-08747; DBLP:conf/emnlp/Li0FT24) is promising. Numerous studies have demonstrated that the initial tokens of a model’s response are key predictors of output safety (Lin2023TheUS; xu2024safedecoding; qi2024safety). Inspired by this, we intuitively propose to preserve the model’s learned safety capabilities by previewing the answer prefixes. The central idea is modifying the training data to preview the answer prefix tokens, thereby reducing the loss on these tokens and minimizing disturbances to the model’s initial tokens. This ensures that the model’s inherent safety mechanisms remain intact, even as it learns new capabilities.
To this end, we introduce LookAhead Tuning, a simple, data-centric, and resource-efficient framework that modifies training data to incorporate answer prefix previews. As shown in Figure 1, LookAhead Tuning achieves the dual objectives of enhancing domain-specific performance while preserving safety alignment. By previewing the initial tokens of the answer, LookAhead Tuning offers a practical and effective solution to mitigating the risks of safety degradation during fine-tuning.
2. Preliminary
2.1. Notation
For clarity, we first introduce the notation used throughout this work. Let $\theta_{0}$ denote the parameters of the pre-trained model, while $\theta$ represents the model parameters being optimized during fine-tuning, and $\theta^{*}$ indicates the optimized parameters after fine-tuning. The dataset is denoted as $D=\{(I_{i},O_{i})\}_{i=1}^{N}$, where $I_{i}$ is the input and $O_{i}$ is the corresponding output for the $i$-th data point, and $N$ is the total number of data points. For each output sequence $O_{i}$, let $n_{i}$ be the number of tokens, $o_{i,t}$ be the $t$-th token, and $o_{i,<t}$ be all tokens preceding position $t$. The loss function is denoted by $\mathcal{L}$. We define $\text{Safe}(\theta)$ as a function that assesses safety of model $\theta$. Additionally, $D_{\text{KL}}$ denotes the Kullback-Leibler (KL) divergence, and $\rho$ represents the correlation coefficient, which quantifies the strength and direction of the linear relationship between variables.
2.2. Vanilla Fine-Tuning
Vanilla Fine-tuning adapts LLMs to downstream tasks by optimizing their parameters to minimize the loss function. For each data point $i$, the loss function $\mathcal{L}$ is typically cross-entropy loss, denoted as:
(1) $\mathcal{L}(O_{i}\mid I_{i},\theta)=-\sum_{t=1}^{n_{i}}\log\mathcal{P}(o_{i,t}\mid I_{i},o_{i,<t},\theta)$
where $\mathcal{P}(\cdot)$ denotes the model’s probability distribution.
2.3. Token-Level Fine-Tuning
To provide finer control over how closely the fine-tuned model follows the pre-trained distribution, we adopt a token-level objective (qi2024safety). Specifically, a per-token coefficient $\beta_{t}\in[0,1]$ determines the degree of alignment at token $t$. $\beta_{t}=0$ corresponds to standard vanilla fine-tuning; and $\beta_{t}=1$ encourages the fine-tuned model to approximate the original pre-trained model’s output distribution at token $t$. The loss function is defined as:
(2) $\mathcal{L}(O_{i}\mid I_{i},\theta)=-\sum_{t=1}^{n_{i}}\max\left(0,\ell_{i,t}\right)$
(3) $\ell_{i,t}=\log\mathcal{P}(o_{i,t}\mid I_{i},o_{i,<t},\theta)-\beta_{t}\log\mathcal{P}(o_{i,t}\mid I_{i},o_{i,<t},\theta_{0})$
2.4. Key Factors in Fine-Tuning Safety
Prior work has identified several key factors influencing model safety during fine-tuning. Specifically, (zhang2024dissecting) and (lin2024unlocking) have demonstrated that fine-tuning and alignment predominantly impact the initial few tokens of model outputs (ji2025first). Furthermore, (xu2024safedecoding) and (qi2024safety) have explored that controlled perturbations of initial tokens can serve as effective defense strategies, fundamentally leveraging the concept of KL divergence. Formally, the difference in safety between the fine-tuned model $\theta$ and the original base model $\theta_{0}$ is negatively correlated ($\rho<0$) with the KL divergence ($D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})$) between their output probability distributions ($\mathcal{P}$) over the first $m$ tokens. This relationship can be formally denoted as:
(4) $\rho\left(\text{Safe}(\theta)-\text{Safe}(\theta_{0}),\ D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})\right)<0$
We also demonstrate this finding in §4.3, showing that greater divergence in the initial tokens corresponds to reduced safety alignment.
Method GSM8K SAMSum Average
RSR JSR UTILITY RSR JSR UTILITY RSR JSR UTILITY
Seed Model 99.39 90.30 26.69 99.39 90.30 25.07 99.39 90.30 25.88
Vanilla FT 96.67 46.97 42.91 69.09 30.61 52.74 82.88 38.79 47.83
SDFT 96.06-0.61 61.21+14.24 34.32-8.59 92.73+23.64 52.73+22.12 30.89-21.85 94.40+11.52 56.97+18.18 32.61-15.22
Constrained SFT 98.79+2.12 50.30+3.33 35.56-7.35 69.70+0.61 36.06+5.45 50.06-2.68 84.25+1.37 43.18+4.39 42.81-5.02
LookAhead Tuning (True) 98.48+1.81 60.61+13.64 38.21-4.70 94.85+25.76 49.39+18.78 51.08-1.66 96.67+13.79 55.00+16.21 44.65-3.18
LookAhead Tuning (Virtual) 99.39+2.72 62.42+15.45 40.79-2.12 96.67+27.58 56.67+26.06 51.69-1.05 98.03+15.15 59.55+20.76 46.24-1.59
Table 1. Main Results. Colored annotations indicate changes relative to the Vanilla FT baseline: red signifies an increase, and green denotes a decrease. Higher values indicate better performance. The best results of each model are marked in bold, and the second-best ones are underlined.
3. Approach: LookAhead Tuning
Our proposed method, LookAhead Tuning, enhances fine-tuning safety through lightweight data modifications rather than architectural changes. As illustrated in Figure 2 with a concrete example, the approach mainly consists of two strategies, True Answer Preview and Virtual Answer Preview, which preview partial answer prefixes during training, and then implement Implicit Token-Level Fine-Tuning and Inference.
3.1. Partial Answer Preview
True Answer Preview.
The True Answer strategy augments the training data by appending the first $m$ tokens of the ground-truth answer to provide explicit prefix guidance. For each training instance $(I_{i},O_{i})$, the input instruction is modified as:
(5) $I_{i}^{\prime}=I_{i}\oplus\text{`` The answer begins with: ''}\oplus O_{i,\leq m}$
where $\oplus$ denotes string concatenation. This modification (solely modifying the instruction) encourages the model to generate the initial tokens correctly, while reducing loss on those tokens. In doing so, it minimizes perturbations in producing the initial output tokens, which is strongly correlated with safety, and thus helps preserve the model’s inherent safety capabilities.
Virtual Answer Preview.
While the True Answer strategy can maintain the model’s safety, it reveals the prefix of the true answers, potentially limiting the model’s ability to learn the complete responses. To mitigate this risk, the Virtual Answer strategy incorporates a prefix $P$, such as “Let’s solve this problem.”, into the response. This ensures that it does not introduce new biases or convey task-specific information. For each training instance $(I_{i},O_{i})$, we perform the following modifications:
(6) $I_{i}^{\prime}=I_{i}\oplus\text{`` The answer begins with: ''}\oplus P$
(7) $O_{i}^{\prime}=P\oplus O_{i}$
This design maintains the safety-preserving benefits of prefix preview while ensuring that the true answer remains unrevealed.
3.2. Token-Level Fine-Tuning & Inference
Implicit Token-Level Fine-Tuning.
The Partial Answer Preview mechanism implicitly realizes an effective token-level fine-tuning scheme (§2.3). By previewing prefixes, the fine-tuned model is encouraged to remain close to the original distribution on early tokens, thereby simulating large $\beta_{t}$ values. For later tokens, no preview is provided, allowing standard fine-tuning with $\beta_{t}=0$. This design naturally aligns the fine-tuned model with the original on initial tokens while optimizing subsequent tokens purely via cross-entropy, achieving a smooth balance between alignment (on the critical initial tokens that determine safety) and utility (on later tokens that drive task-specific learning). As a result, LookAhead Tuning balances the dual objectives of capability acquisition and safety preservation in a simple, data-driven manner.
Inference.
At inference time, LookAhead Tuning requires no modification to the input data. The model behaves like a standard fine-tuned LLM, ensuring practical deployment. To guarantee reproducibility, we employ greedy decoding for response generation. Given an input $I$, the model generates the output sequence $O^{*}=(o_{1}^{*},\ldots,o_{T}^{*})$ by iteratively selecting the highest-probability token from the vocabulary $V$, denoted as:
(8) $o_{t}^{*}=\arg\max_{o\in V}\mathcal{P}(o\mid I,o_{<t},\theta^{*})$
4. Experiments
4.1. Setup
We fine-tune LLaMA2-7B-Chat (Touvron2023Llama2O) on GSM8K (cobbe2021training) and SAMSum (samsum) datasets for utility and safety evaluation. For our LookAhead Tuning method, we set the number of previewed tokens $m=6$ for the True Answer Preview strategy and use the buffer prefix $P=\textit{``Let's solve this problem.''}$ for the Virtual Answer Preview strategy. We compare our method against baselines including the Seed Model, Vanilla Fine-Tuning (FT), SDFT (DBLP:conf/acl/YangPFWCZL24), and Constrained SFT (qi2024safety). Utility is measured by Accuracy for GSM8K and ROUGE-1 for SAMSum. Safety is evaluated with the HEx-PHI dataset (qi2023fine) under the protocol of (qi2024safety), where Raw Safe Rate (RSR) denotes the proportion of safe responses to direct harmful queries, and Jailbreak Safe Rate (JSR) denotes the proportion of safe responses under adversarial prefilling attacks. We fine-tune the model for 3 epochs with a learning rate of $2\times 10^{-5}$ and the batch size of 64, using AdamW (betas 0.5, 0.999). Training has been conducted on 4×NVIDIA A100 (80GB) with bf16 precision and DeepSpeed ZeRO Stage 2 for efficiency.
4.2. Main Results
As shown in Table 1, both strategy variants of LookAhead Tuning (True/Virtual Answer Preview) perform well across safety and utility evaluations compared to baselines. The True Answer Preview Strategy, i.e., LookAhead Tuning (True), achieves the second-best performance in terms of RSR and Utility by solely modifying the instruction with explicit prefix guidance. The Virtual Answer Preview Strategy, i.e., LookAhead Tuning (Virtual), attains state-of-the-art performance by jointly modifying instruction and answer with synthetic prefixes, ranking first across all metrics. These findings highlight the effectiveness of prefix-based previews for maintaining model safety while enabling task adaptation. In addition to strong performance, LookAhead Tuning remains resource-efficient. As detailed in Table 2, our true and virtual answer preview strategies respectively increase FLOPs by only $2.18\%$ and $3.90\%$ compared to Vanilla Fine-Tuning, demonstrating the resource efficiency of our approach. Overall, the true answer preview strategy is particularly suited for low-resource scenarios where revealing a small number of answer tokens is acceptable, while the virtual answer preview strategy offers stronger safety guarantees without exposing true prefixes, although slightly increasing resource consumption. Together, these two strategy variants, by jointly modifying instructions and answers with prefixes, provide flexible mechanisms that balance safety preservation and task performance.
Method GSM8K SAMSum Average
Vanilla Fine-Tuning $3.38\times 10^{17}$ $9.43\times 10^{17}$ $6.41\times 10^{17}$
True Answer Preview $3.47\times 10^{17}$ $9.62\times 10^{17}$ $6.55\times 10^{17}$
Virtual Answer Preview $3.55\times 10^{17}$ $9.76\times 10^{17}$ $6.66\times 10^{17}$
Table 2. Average FLOPs of LookAhead Tuning (True & Virtual Answer Preview) and Vanilla Fine-Tuning. Our method introduces marginal overhead compared to Vanilla Fine-Tuning.
4.3. Further Analysis
Fine-tuning safety is closely related to KL divergence of early tokens.
To better understand this relationship, we analyze the per-token KL divergence of first few tokens in responses to harmful prompts, between the original and fine-tuned models on the harmful HEX-PHI dataset (qi2023fine; qi2024safety). As shown in Figure 3, our method significantly reduces KL divergence for first few (2 & 4) tokens compared to vanilla fine-tuning. In particular, KL divergence values for 6 and 8 tokens remain similar to baseline models, indicating that LookAhead Tuning primarily constrains earliest tokens while allowing later tokens to adapt freely. This distribution confirms that our method minimizes perturbations to critical prefix tokens, thereby enhancing safety. These results align with the theoretical framework introduced in §2.4, which predicts a negative correlation between safety retention and KL divergence of early tokens. In practice, we observe that preserving low KL divergence on initial tokens ($D^{(m)}_{\text{KL}}(\mathcal{P}_{\theta}\|\mathcal{P}_{\theta_{0}})\downarrow$) directly corresponds to stronger safety alignment ($\text{Safe}(\theta)\uparrow$) without hindering downstream performance.
Figure 3. Per-token KL Divergence between the fine-tuned models and the original model on harmful HEx-PHI dataset.
The more true answer tokens, the safer the model, but at the cost of downstream task performance.
As illustrated in Figure 4, increasing the number of previewed true tokens consistently improves model safety, with larger RSR and JSR. However, this gain comes with a trade-off: performance on downstream tasks tends to decline as more tokens are revealed, with smaller UTILITY. This pattern highlights an important design consideration: While previewing additional tokens strengthens safety alignment, excessive exposure can hinder the model’s ability to adapt effectively to task-specific learning. Striking the right balance in the number of previewed tokens is therefore essential for maximizing safety benefits while preserving downstream utility.
Figure 4. Effect of varying the number of previewed true answer tokens on model safety and utility.
Prefix variations preserve robustness in the virtual answer preview strategy.
To evaluate the robustness of the virtual answer strategy, we experimented with two alternative prefixes: a semantically coherent phrase (“Let’s deal with this situation.”) and a nonsensical string (“dadjalsasdfghkjzmnb”). As shown in Figure 5, the coherent phrase achieves performance comparable to the default template, while the nonsensical string produces only a minor decline, attributable to its lack of semantic structure. Importantly, in both cases the model maintains strong safety alignment and competitive task performance, demonstrating that the virtual answer preview strategy is resilient to prefix variations and does not rely on carefully crafted phrasing.
Figure 5. Effectiveness of different prefixes using the virtual answer preview strategy. p1: “Let’s solve this problem.”; p2: “Let’s deal with this situation.”; p3: “dadjalsasdfghkjzmnb”.
5. Conclusion and Future Work
In this work, we introduced LookAhead Tuning, a simple, data-centric, and resource-efficient approach to fine-tuning LLMs that balances task performance and safety. By previewing partial prefixes, either true or virtual, our method preserves the model’s inherent safety mechanisms while achieving competitive downstream task performance. Extensive experiments confirm its effectiveness and efficiency across multiple benchmarks.
Looking forward, we plan to extend LookAhead Tuning in two directions. First, we aim to explore broader architectures, including multimodal LLMs, and test robustness under diverse adversarial settings. Second, we will investigate automated strategies for optimizing the trade-off between task utility and safety alignment, potentially adapting the preview length or prefix type dynamically.
Acknowledgments
This work was supported by the Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, Ant Group and Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph, Ningbo Natural Science Foundation (2024J020), and the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00).
Ethics Statement
This work adheres to ethical AI research guidelines, striving to ensure safety in model fine-tuning. Although our approach has proven effective in reducing harmful output, we recognize the need to continuously evaluate edge cases and unexpected behaviors in real-world deployment. Moreover, we categorically denounce any malicious misuse of this technology, striving to ensure that the development of AI consistently aligns with human ethical principles.